OpenStudy (anonymous):

Create an exponential growth function, f(x), to model a population of frogs that is growing every year. Identify the principal amount, the growth rate, and the appropriate domain and range for your function. Explain how these key features would affect the graph of f(x). Using your function f(x) from question 2. Demonstrate and explain how to find the average rate of change between year 3 and 5, and between year 5 and 7. Explain what the average rate of change represents to the frog population. My function is, F(x) =150(1.02)^6

OpenStudy (anonymous):

just need help with number 2

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (the_fizicx99):

Hmm, well the principle or P is 150 in your function, the domain would be the x that doubles every year. The y intercept is also the principle, the growth rate is 1.02. Those are your key features.

OpenStudy (the_fizicx99):

So in your function, F(x) =150(1.02)^6 there shouldn't be a 6, it should only be F(x) =150(1.02)^x

OpenStudy (the_fizicx99):

x will be any year you want it to be, they're asking for 3 and 5, and 5 & 7

OpenStudy (the_fizicx99):

F(x) =150(1.02)^x F(x) =150(1.02)^3 F(3) =150(1.02^3) f(3) = 150(1.061208) f(3) = 159.1812, so just use 159 as there can't be a decimal we're taking about living things, there can't be half a frog lol >.< unless its dead though >.> f(3) = 159 You do f(5) and f(7), I'll check them over

OpenStudy (anonymous):

Year 5, F(5) = 150(1.02)^5 166 Year 7, F(7) = 150(1.02)^7 172

OpenStudy (the_fizicx99):

Good, now the average rate of change formula is; \[\frac{ f(b) - f(a) }{b -a}\]

OpenStudy (the_fizicx99):

Because its exponential their average rate of change will be different

OpenStudy (the_fizicx99):

What did you get